NvidiaIntely (nvidiaintely):

Find the number of moles in 60 grams of SO3. 0.039 moles 0.75 moles 1.25 moles 1.33 moles

NvidiaIntely (nvidiaintely):

@welshfella @Will.H @TheSmartOne @Kevin @brainlesslonelygirl

OpenStudy (welshfella):

1 mole of SO3 = 32 + 3*16

OpenStudy (welshfella):

So 60 g = 60 / 80 g

OpenStudy (kevin):

mol = gr / MR = 60/(32 + (3*16)) = 0.75

OpenStudy (welshfella):

* 60/80 moles = 0.75 moles

NvidiaIntely (nvidiaintely):

Thanks Welshfella and Kevin!

OpenStudy (kevin):

u're welcome :)